Mr. and Mrs. J. B. Davis just moved into a new home and need to purchase kitchen appliances. The Davises are going to purchase a refrigerator, a stove, and a dishwasher. They have a number of choices for each appliance.
For Refrigerator: General Electric, Maytag, Kenmore, Acme For Stove: General Electric, Frigidaire, Roper, Acme For Dishwasher: General Electric, KitchenAid, WhirlPool
Find the probability that they select no General Electric appliances.



Answer by stanbon(75887)

Mr. and Mrs. J. B. Davis just moved into a new home and need to purchase kitchen appliances. The Davises are going to purchase a refrigerator, a stove, and a dishwasher. They have a number of choices for each appliance.
For Refrigerator: General Electric, Maytag, Kenmore, Acme
For Stove: General Electric, Frigidaire, Roper, Acme
For Dishwasher: General Electric, KitchenAid, WhirlPool
Find the probability that they select no General Electric appliances.
---------------------------
Prob(no GE refrig) = 3/4
Prob(no GE stove) = 3/4
Prob(no GE dishw) = 2/3
-----------------------------------
Prob(none of the above) = (3/4)^2 * (2/3) = 18/48 = 3/8
==============================
Cheers,
Stan H.